OpenStudy (australopithecus):

A 15.0-kg block is dragged over a rough, horizontal surface by a 70.0-N force acting at 20.0° above the horizontal. The block is displaced 5.00 m, and the coefficient of kinetic friction is 0.300. Find the work done on the block by c) the gravitational force. (d) What is the increase in internal energy of the block–surface system owing to friction? (e) Find the total change in the block’s kinetic energy.

OpenStudy (australopithecus):

I assume the free body diagram is

OpenStudy (australopithecus):

frictional force = 44.145N

OpenStudy (australopithecus):

I messed up my free body diagram Normal force 15(0) = N + 70sin(20) - 15(9.81) N = -70sin(20) + 15(9.81) Normal Force = 123N

OpenStudy (australopithecus):

Nevermind I solved it :)

OpenStudy (australopithecus):

My frictional force was wrong :S but I corrected it when I corrected the normal froce I had d) 184N e)

OpenStudy (anonymous):

that's great! :)

OpenStudy (anonymous):

what did u do to find (d) ?

OpenStudy (australopithecus):

I assume that the graviational and normal force both equal to 0 work, for d I got 144N m

OpenStudy (australopithecus):

or 144J

OpenStudy (australopithecus):

I messed up my units oh well

OpenStudy (anonymous):

yes, gravitational force does 0 work...since there is no disp in that direction. and the answer to (e) is just the work done in displacing the block by 5m. but i don't know how to solve for (d)...can u help me there?

OpenStudy (australopithecus):

Yes not a problem To find the total change we use equation K = 1/2mv^2 we need to find velocity so we use the equation vf = vi + 2a(Delta x) thus we have vf = 0 + 2a(5) now we need to solve for acceleration we can jsut use the force equation we know that ma = 70cos(20) - 36.9 = 15a = 70cos(20) - 36.9 the it is as simple as solving for a and then subing vf into the kinetic energy equation

OpenStudy (australopithecus):

wow internet explorer is horriable, no idea why my school doesnt have fire fox

OpenStudy (anonymous):

no...i meant the how to find the increase in the internal energy of the block-surface system because of friction..?

OpenStudy (anonymous):

side note: u don't use google chrome?

OpenStudy (australopithecus):

no I prefer firefox. I found the Normal force to be 123N so then I use that to find the frictional force which is 36.9N then I jus sub the frictional force into the work equation W = 36.9N * 5 = 184J

OpenStudy (australopithecus):

hope that helps I have to leave Im going to be late for my dgd, thanks for confirming for me :)

OpenStudy (anonymous):

oh got it! so that's what it means! thanks! and how come ur friction force came out to be in 30s? Friction force = u*N = 3*123 = 369 N

OpenStudy (australopithecus):

u = 0.3

OpenStudy (anonymous):

sry!!!! its 0.3! just saw that! my bad!

OpenStudy (australopithecus):

no problem thanks again

OpenStudy (anonymous):

u too!